1) (y - xz)p + (yz - x)q = (x + y)(x - y)


1
Expert's answer
2021-06-28T18:04:11-0400

Given equation is "(y - xz)p + (yz - x)q = (x + y)(x - y)"


Comparing with the standard equation, "Pp+Qq=R"


"\frac{dx}{P}=\frac{dy}{Q}=\frac{dz}{R}"


Equation will be,

"\frac{dx}{y-xz}=\frac{dy}{yz-x}=\frac{dz}{(x+y)(x-y)} = \frac{xdx+ydy+zdz}{0} = \frac{ydx+xdy}{y^2-x^2}"


"\frac{dx}{y-xz}=\frac{dy}{yz-x}=\frac{dz}{x^2-y^2} = \frac{xdx+ydy+zdz}{0} = \frac{ydx+xdy}{y^2-x^2}"


"{xdx+ydy+zdz} = {0}"

solving it,

"x^2+y^2+z^2=c_1" (1)


Equation (1) is the solution of the given differential equation


Taking (iii) and last part,

"\frac{dz}{x^2-y^2} = \frac{ydx+xdy}{y^2-x^2}"


"dz = -ydx-xdy"


Integrating both sides,

"z = -xy + c_2"  (2)


"z +xy = c_2"


Equations (1) and (2) are solutions to the equation.


The general solution of the differential equation will be

"\phi(x^2+y^2+z^2,z+xy)=0."